Kylie9400:

I need help I dont know how to do this problem 4/5 + 3/4

luvinmycoke:

1. Find a common denominator. 16/20 + 15/20 2. Add 16+15/20 = 31/ 20 3. Simplify 31/20 = 1 and 11/20